zarkam21 (zarkam21):

If r is the radius of a circle and d is its diameter, which of the following is an equivalent formula for the circumference C=2πr? A. C=2πd B. C=πD C. C=πr2 D. C=πdr

rishavraj (rishavraj):

see diameter = 2*radius and circumference = 2*pi*r r -> radius

rishavraj (rishavraj):

so u can write 2r = d

zarkam21 (zarkam21):

But thats not even a choice

OpenStudy (michele_laino):

hint: \[C = 2\pi r = \pi \left( {2r} \right) = ...?\]

rishavraj (rishavraj):

see C = 2*pi*r as I told r = d/2 so plug tht in the equation :) u will get c = pi*d

zarkam21 (zarkam21):

Choice B?

rishavraj (rishavraj):

yup :))

zarkam21 (zarkam21):

OKay thanks